# MMRAG-DocQA: A Multi-Modal Retrieval-Augmented Generation Method for Document Question-Answering with Hierarchical Index and Multi-Granularity Retrieval

### v2

@@ Abstract
The multi-modal long-context document question-answering task aims to locate and integrate multi-modal evidences (such as texts, tables, charts, images, and layouts) distributed across multiple pages, for question understanding and answer generation. The existing methods can be categorized into Large Vision-Language Model (LVLM)-based and Retrieval-Augmented Generation (RAG)-based methods. However, the former were susceptible to hallucinations, while the latter struggled for inter-modal disconnection and cross-page fragmentation. To address these challenges, a novel multi-modal RAG model, named MMRAG-DocQA, was proposed, leveraging both textual and visual information across long-range pages to facilitate accurate question answering. A hierarchical indexing method with the integration of flattened in-page chunks and topological cross-page chunks was designed to jointly establish in-page multi-modal associations and long-distance cross-page dependencies. By means of joint similarity evaluation and large language model (LLM)-based re-ranking, a multi-granularity semantic retrieval method, including the page-level parent page retrieval and document-level summary retrieval, was proposed to foster multi-modal evidence connection and long-distance evidence integration and reasoning. Experimental results performed on public datasets, MMLongBench-Doc and LongDocURL, demonstrated the superiority of our MMRAG-DocQA method in understanding and answering modality-rich and multi-page documents. Our code is available at https://github.com/Gzy1112/MMRAG-DocQA.
@@ Introduction
Considering the above challenges and solution ideas, a novel retrieval-augmented generation method, named MMRAG-DocQA, was proposed for the multi-modal long-context document question-answering task. The major contributions can be summarized as follows:
  • In proposed MMRAG-DocQA, a hierarchical index structure with flattened in-page and topological cross-page chunks was proposed to establish correlations between diverse modalities and multiple pages.
@@
Figure 2: Overview of MMRAG-DocQA with hierarchical index and multi-granularity retrieval for multi-modal Doc-QA. ($P_{n}$ is the parent page of $c_{n}$, $cls_{K}$ is the clustered block, ‘CoT’ denotes chain-of-thought and ‘SO’ denotes a structured output format.)
Our MMRAG-DocQA Methodology
@@
Figure 2 presents an overview of our MMRAG-DocQA method. Given a query $Q$ and its corresponding document $D=\{d_{1},d_{2},...,d_{N}\}$, where $N$ is denoted as the page number and $d_{i}$ represents the $i$-th page.The multi-modal long-context Doc-QA can be formalized as follows:
@@
MMRAG-DocQA (OURS)(page=4) 41.6 30.2 40.9 48.9 25.1 48.5 31.7 74.9 48.2 41.4
MMRAG-DocQA (OURS)(page=10) 45.9 34.4 44.9 51.1 37.5 53.5 36.8 76.2 52.3 46.0
@@
MMRAG-DocQA(OURS)(page=4) 52.4
MMRAG-DocQA(OURS)(page=10) 57.2
@@
We compared our MMRAG-DocQA with existing SOTA LVLM/LLM-based and RAG-based methods on the MMLongBench-Doc and LongDocURL datasets.
MMLongBench-Doc. Table 1 listed the performance of models on the MMLongBench-Doc dataset. We observed that: (1) Both LVLM-based methods and LLM-based methods with OCR-parsed documents exhibited poor performance and struggled with multi-modal comprehension and long-distance reasoning for long-context document. Our MMRAG-DocQA surpassed the best-performing LVLM, i.e., GPT-4V, by 19.9% and 14.8% in generalized accuracy and F1 score, respectively. (2) Compared with the RAG-based SOTA M3DocRAG, when the page number was four, MMRAG-DocQA achieved superiority on all metrics, with improvements of 27.2% and 18.8% on generalized accuracy and F1 score. When the page number was set to ten, MMRAG-DocQA achieved better performance with an accuracy of 52.3% and a F1 score of 46.0%. To be specific, the accuracy of visual charts (CHA) and figures (FIG) separately increased to 44.9% and 37.5%. The accuracy of multi-page (MUL) and unanswerable-questions (UNA) attained 36.8% and 76.2%, respectively. (3) These experimental results proved the advantages of our model in multiple modality understanding and long-distance reasoning, and the capability of alleviating hallucinations caused by LLMs.
LongDocURL. Table 2 showed the performance of models on the LongDocURL dataset. Similar phenomena were found: (1) Our MMRAG-DocQA achieved better performance, compared to the top-performing LVLM (i.e., Qwen2-VL) and LLM (i.e., o1-preview). (2) MMRAG-DocQA method surpassed the current SOTA M3DocRAG by a margin of 5% in terms of generalized accuracy, when the page number was set to ten. (3) These results further validated the effectiveness of our approach for the connection of multi-modality and the link of cross-page evidence.
@@
Figure 3: The trend of our MMRAG-DocQA model performance changing with the page number and summary number on the MMLongBench-Doc dataset.
@@
Figure 3 showed the parameter analysis of retrieved page number and summary number for answer generation. MMRAG-DocQA reached peak accuracy and F1 score when the parent page count was set to 10, but began to decline when the number of pages exceeded 10, which suggested that blindly increasing the page count may negatively impact performance. Meanwhile, as the number of summaries increased, the performance of MMRAG-DocQA initially improved and subsequently declined, reaching its peak when the summary count was 10, which further supported the notion that an excessive amount of content may introduce irrelevant or distracting information, thereby negatively impacting the quality of the responses. Therefore, we set the number of pages and summaries to 10 to ensure that the retrieved content contains more evidence and avoids noisy information.
@@
Figure 4: Case Study on dataset MMLongBench-Doc and LongDocURL to compare the answer response of our MMRAG-DocQA and LVLM-based methods (such as GPT-4o, DeepSeek-chat, Qwen-VL-Plus and ERNIE-Turbo).
@@
As shown in Table 5, four different LLMs were selected as the cornerstone models to evaluate the performance impact on the document question-answering task. We found that Qwen-turbo achieved the highest accuracy at 52.3% on the MMLongBench-Doc dataset, while GPT-4o attained the best performance with an accuracy of 57.2% on the LongDocURL dataset. These results demonstrated that our proposed Retrieval-Augmented Generation (RAG)-based method, MMRAG-DocQA, achieved excellent response on different large language models, which again validated the universality and transferability of our method.
@@
Figure 4 demonstrated two cases from the MMLongBench-Doc and LongDocURL datasets. The first question asked the comparison of fan numbers, with the answer required integrating textual statistics of Vietnam from page 28 with visual data of Appota on page 31. VLVM-based methods, such as Deepseek-chat and Qwen-VL-Plus, failed to answer the question. However, MMRAG-DocQA correctly retrieved fan count from separate pages and concluded that Appota had more followers. The second question asked the number of 1492-MCA branch circuit breakers’ rating, with the answer found in structured tables across page 110 and 111. GPT-4o and Qwen-VL-Plus gave incorrect answers of 9 and 10, respectively, which failed to identify all the entries. However, MMRAG-DocQA correctly located the relevant table and identified the rating values. In summary, these cases again demonstrated that our method is adaptable for the comprehension of multi-modal and multi-page evidences that retrieved from document content.
@@ Conclusion
A retrieval-augmented generation method (MMRAG-DocQA) was presented for multi-modal long-context document question-answering. A hierarchical index structure with flattened in-page and topological cross-page index was constructed to establish multi-modal connection and long-distance linkage. A multi-granularity retrieval with page-level parent page retrieval and document-level summary retrieval was proposed for searching required evidences, which were scattered across multi-modalities and multi-pages. Experiments conducted on two public datasets demonstrated the superiority of MMRAG-DocQA in multi-modal long-context Doc-QA.
@@ Acknowledgments
Figure 5 illustrates the capabilities required for multi-modal long-context Doc-QA and the advantage of our proposed MMRAG-DocQA method. For the question “What percentage of respondents of the sector in which 15% are doing promotions to customers over Wi-Fi use wifi at stores?”, evidences required for the answer are scattered in the visual charts on page 11 and page 14. To correctly answer this question, multi-modal Doc-QA methods need to establish multi-modal connections. Due to the common keywords and similar semantics, the question is relatively easy to retrieve the textual titles “% RESPONDENTS USING WIFI AT STORES” in page 11 and “Are you doing promotions to customers over Wi-Fi?” in page 14. Therefore, it is crucial to associate these textual captions with their surrounding related visual bar charts. Multi-modal Doc-QA methods also need to achieve long-distance reasoning to synthesize information on these non-consecutive pages. To be specific, Doc-QA methods first require to identify the “Hospitality” industry (the only category with a 15% promotion from the visual chart on page 14), and then retrieve its corresponding 100% Wi-Fi usage rate from the data on page 11.
Our MMRAG-DocQA first confirmed from page 14 that hospitality is the only industry that conducts a 15% promotion through Wi-Fi, then found the proportion of respondents in the hotel industry who use Wi-Fi in stores from page 11, and finally concluded the correct answer 100 through reasonable multi-step reasoning. This indicates that our MMRAG-DocQA has the ability of modality correlation and long-distance reasoning, which is crucial for integrating multi-modal cross-page evidences to answer questions.
@@
In the settings of our MMRAG-DocQA, we used docling to achieve pdf parsing and used off-the-shelf Large Vision-Language Models (such as Qwen-vl-plus) to describe visual information in documents. For the flattened in-page index, the chunk size was set to 300 tokens (about 15 sentences) and encoded by text-embedding-v4 provided by Qwen. To avoid information loss due to text split, a text overlap with 50 tokens was added. For the topological cross-page index, the chunk size was set to 100 (about 5 sentences) and encoded by multi-qa-mpnet-base-cos-v1 provided by Sentence-Transformer. The gaussian mixture model was used for chunk clustering, and Gpt-3-turbo was adopted for summarizing the clustered chunks. For multi-granularity evidence retrieval, we selected 10 highest-scored parent pages and summaries as the input context for each question. We used Large Vision-Language Models (such as Qwen-turbo) to conduct retrieval re-ranking. We directly used off-the-shelf LLMs for answer generation. All experiments were conducted on a single NVIDIA A100 GPU.

### v3

# MHier-RAG: Multi-Modal RAG for Visual-Rich Document Question-Answering via Hierarchical and Multi-Granularity Reasoning

@@ Abstract
The multi-modal long-context document question-answering task aims to locate and integrate multi-modal evidences (such as texts, tables, charts, images, and layouts) distributed across multiple pages, for question understanding and answer generation. The existing methods can be categorized into Large Vision-Language Model (LVLM)-based and Retrieval-Augmented Generation (RAG)-based methods. However, the former were susceptible to hallucinations, while the latter struggled for inter-modal disconnection and cross-page fragmentation. To address these challenges, a novel multi-modal RAG model, named MHier-RAG, was proposed, leveraging both textual and visual information across long-range pages to facilitate accurate question answering for visual-rich documents. A hierarchical indexing method with the integration of flattened in-page chunks and topological cross-page chunks was designed to jointly establish in-page multi-modal associations and long-distance cross-page dependencies. By means of joint similarity evaluation and large language model (LLM)-based re-ranking, a multi-granularity semantic retrieval method, including the page-level parent page retrieval and document-level summary retrieval, was proposed to foster multi-modal evidence connection and long-distance evidence integration and reasoning. Experimental results performed on public datasets, MMLongBench-Doc and LongDocURL, demonstrated the superiority of our MHier-RAG method in understanding and answering modality-rich and multi-page documents.
@@ Introduction
Considering the above challenges and solution ideas, a novel retrieval-augmented generation method, named MHier-RAG, was proposed for the multi-modal long-context document question-answering task. The major contributions can be summarized as follows:
  • In proposed MHier-RAG, a hierarchical index structure with flattened in-page and topological cross-page chunks was proposed to establish correlations between diverse modalities and multiple pages.
@@
Figure 2: Overview of MHier-RAG with hierarchical index and multi-granularity retrieval for multi-modal Doc-QA. ($P_{n}$ is the parent page of $c_{n}$, $cls_{K}$ is the clustered block, ‘CoT’ denotes chain-of-thought and ‘SO’ denotes a structured output format.)
Our MHier-RAG Methodology
@@
Figure 2 presents an overview of our MHier-RAG method. Given a query $Q$ and its corresponding document $D=\{d_{1},d_{2},...,d_{N}\}$, where $N$ is denoted as the page number and $d_{i}$ represents the $i$-th page.The multi-modal long-context Doc-QA can be formalized as follows:
@@
MHier-RAG (OURS)(page=4) 41.6 30.2 40.9 48.9 25.1 48.5 31.7 74.9 48.2 41.4
MHier-RAG (OURS)(page=10) 45.9 34.4 44.9 51.1 37.5 53.5 36.8 76.2 52.3 46.0
@@
MHier-RAG(OURS)(page=4) 52.4
MHier-RAG(OURS)(page=10) 57.2
@@
We compared our MHier-RAG with existing SOTA LVLM/LLM-based and RAG-based methods on the MMLongBench-Doc and LongDocURL datasets.
MMLongBench-Doc. Table 1 listed the performance of models on the MMLongBench-Doc dataset. We observed that: (1) Both LVLM-based methods and LLM-based methods with OCR-parsed documents exhibited poor performance and struggled with multi-modal comprehension and long-distance reasoning for long-context document. Our MHier-RAG surpassed the best-performing LVLM, i.e., GPT-4V, by 19.9% and 14.8% in generalized accuracy and F1 score, respectively. (2) Compared with the RAG-based SOTA M3DocRAG, when the page number was four, MHier-RAG achieved superiority on all metrics, with improvements of 27.2% and 18.8% on generalized accuracy and F1 score. When the page number was set to ten, MHier-RAG achieved better performance with an accuracy of 52.3% and a F1 score of 46.0%. To be specific, the accuracy of visual charts (CHA) and figures (FIG) separately increased to 44.9% and 37.5%. The accuracy of multi-page (MUL) and unanswerable-questions (UNA) attained 36.8% and 76.2%, respectively. (3) These experimental results proved the advantages of our model in multiple modality understanding and long-distance reasoning, and the capability of alleviating hallucinations caused by LLMs.
LongDocURL. Table 2 showed the performance of models on the LongDocURL dataset. Similar phenomena were found: (1) Our MHier-RAG achieved better performance, compared to the top-performing LVLM (i.e., Qwen2-VL) and LLM (i.e., o1-preview). (2) MHier-RAG method surpassed the current SOTA M3DocRAG by a margin of 5% in terms of generalized accuracy, when the page number was set to ten. (3) These results further validated the effectiveness of our approach for the connection of multi-modality and the link of cross-page evidence.
@@
Figure 3: The trend of our MHier-RAG model performance changing with the page number and summary number on the MMLongBench-Doc dataset.
@@
Figure 3 showed the parameter analysis of retrieved page number and summary number for answer generation. MHier-RAG reached peak accuracy and F1 score when the parent page count was set to 10, but began to decline when the number of pages exceeded 10, which suggested that blindly increasing the page count may negatively impact performance. Meanwhile, as the number of summaries increased, the performance of MHier-RAG initially improved and subsequently declined, reaching its peak when the summary count was 10, which further supported the notion that an excessive amount of content may introduce irrelevant or distracting information, thereby negatively impacting the quality of the responses. Therefore, we set the number of pages and summaries to 10 to ensure that the retrieved content contains more evidence and avoids noisy information.
@@
Figure 4: Case Study on dataset MMLongBench-Doc and LongDocURL to compare the answer response of our MHier-RAG and LVLM-based methods (such as GPT-4o, DeepSeek-chat, Qwen-VL-Plus and ERNIE-Turbo).
@@
As shown in Table 5, four different LLMs were selected as the cornerstone models to evaluate the performance impact on the document question-answering task. We found that Qwen-turbo achieved the highest accuracy at 52.3% on the MMLongBench-Doc dataset, while GPT-4o attained the best performance with an accuracy of 57.2% on the LongDocURL dataset. These results demonstrated that our proposed Retrieval-Augmented Generation (RAG)-based method, MHier-RAG, achieved excellent response on different large language models, which again validated the universality and transferability of our method.
@@
Figure 4 demonstrated two cases from the MMLongBench-Doc and LongDocURL datasets. The first question asked the comparison of fan numbers, with the answer required integrating textual statistics of Vietnam from page 28 with visual data of Appota on page 31. VLVM-based methods, such as Deepseek-chat and Qwen-VL-Plus, failed to answer the question. However, MHier-RAG correctly retrieved fan count from separate pages and concluded that Appota had more followers. The second question asked the number of 1492-MCA branch circuit breakers’ rating, with the answer found in structured tables across page 110 and 111. GPT-4o and Qwen-VL-Plus gave incorrect answers of 9 and 10, respectively, which failed to identify all the entries. However, MHier-RAG correctly located the relevant table and identified the rating values. In summary, these cases again demonstrated that our method is adaptable for the comprehension of multi-modal and multi-page evidences that retrieved from document content.
@@ Conclusion
A retrieval-augmented generation method (MHier-RAG) was presented for multi-modal long-context document question-answering. A hierarchical index structure with flattened in-page and topological cross-page index was constructed to establish multi-modal connection and long-distance linkage. A multi-granularity retrieval with page-level parent page retrieval and document-level summary retrieval was proposed for searching required evidences, which were scattered across multi-modalities and multi-pages. Experiments conducted on two public datasets demonstrated the superiority of MHier-RAG in multi-modal long-context Doc-QA.
@@ Acknowledgments
Figure 5 illustrates the capabilities required for multi-modal long-context Doc-QA and the advantage of our proposed MHier-RAG method. For the question “What percentage of respondents of the sector in which 15% are doing promotions to customers over Wi-Fi use wifi at stores?”, evidences required for the answer are scattered in the visual charts on page 11 and page 14. To correctly answer this question, multi-modal Doc-QA methods need to establish multi-modal connections. Due to the common keywords and similar semantics, the question is relatively easy to retrieve the textual titles “% RESPONDENTS USING WIFI AT STORES” in page 11 and “Are you doing promotions to customers over Wi-Fi?” in page 14. Therefore, it is crucial to associate these textual captions with their surrounding related visual bar charts. Multi-modal Doc-QA methods also need to achieve long-distance reasoning to synthesize information on these non-consecutive pages. To be specific, Doc-QA methods first require to identify the “Hospitality” industry (the only category with a 15% promotion from the visual chart on page 14), and then retrieve its corresponding 100% Wi-Fi usage rate from the data on page 11.
Our MHier-RAG first confirmed from page 14 that hospitality is the only industry that conducts a 15% promotion through Wi-Fi, then found the proportion of respondents in the hotel industry who use Wi-Fi in stores from page 11, and finally concluded the correct answer 100 through reasonable multi-step reasoning. This indicates that our MHier-RAG has the ability of modality correlation and long-distance reasoning, which is crucial for integrating multi-modal cross-page evidences to answer questions.
@@
In the settings of our MHier-RAG, we used docling to achieve pdf parsing and used off-the-shelf Large Vision-Language Models (such as Qwen-vl-plus) to describe visual information in documents. For the flattened in-page index, the chunk size was set to 300 tokens (about 15 sentences) and encoded by text-embedding-v4 provided by Qwen. To avoid information loss due to text split, a text overlap with 50 tokens was added. For the topological cross-page index, the chunk size was set to 100 (about 5 sentences) and encoded by multi-qa-mpnet-base-cos-v1 provided by Sentence-Transformer. The gaussian mixture model was used for chunk clustering, and Gpt-3-turbo was adopted for summarizing the clustered chunks. For multi-granularity evidence retrieval, we selected 10 highest-scored parent pages and summaries as the input context for each question. We used Large Vision-Language Models (such as Qwen-turbo) to conduct retrieval re-ranking. We directly used off-the-shelf LLMs for answer generation. All experiments were conducted on a single NVIDIA A100 GPU.